how many possible telephone numbers are there when there are seven digits the first two of which between 2 and 9 inclusive the third digit between 1 and 9 inclusive and each of the remaining may be between 0 and 9 inclusive ?



1
Expert's answer
2022-02-03T18:13:30-0500

There are 8 possabilities for each of the first two numbers, 9 possabilities for the third number and 10 possabilities for each of the remaining four numbers. So, there are totally

"8*8*9*10*10*10*10=8^2*9*10^4=5760000" phone numbers